Pat is required to sell candy bars until five of them have been sold. So the child goes door to door. At each house, there is 0.4 probability to sell a candy bar. a) what is probability that Pat finishes on seventh house or less? b) what is the average number of houses Pat should visit to sell 3 candy bars
Answer #1

probability that Pat finishes on seventh house or less=1-P(sell at most 4 candy till 7th house)

=1-P(X<=4)

=1-(7C0)*(0.4)0(0.6)7-(7C1)*(0.4)1(0.6)6-(7C2)*(0.4)2(0.6)5-(7C3)*(0.4)3(0.6)4-(7C4)*(0.4)4(0.6)3

=0.0963

b)

average number of houses Pat should visit to sell 3 candy bars =r/p=5/0.4 =12.5